# Unfitted finite element modelling of surface-bulk viscous flows in animal cells

### v1

@@ Abstract.
CIRB, Collège de France, Université PSL, CNRS, INSERM, 75005 Paris, France.
@@
E-mails: eric.neiva@college-de-france.fr (EN), herve.turlier@college-de-france.fr (HT)
@@ 1. Introduction
Morphogenesis is the biological process by which a cell, tissue or organism takes shape and develops its distinctive form [lecuit2007cell]. Researchers in this field seek to understand the underlying genetic, cellular, mechanical, and environmental factors that drive the formation of complex biological structures. The ultimate goal is to unravel the fundamental principles that govern the development of living organisms. Understanding morphogenesis has significant far-reaching implications for various fields, including regenerative medicine [sasai2013next], cancer research [yamada2007modeling] or environmental and ecological studies [wichard2015green].
@@
This work centres upon the morphogenesis of single animal cells. Animal cells maintain their shape primarily through an effective surface tension, adopting shapes akin to soap bubbles [hayashi2004surface]. The main contributor to this effective surface tension in animal cells is the actomyosin cortex [clark2014stresses]: a thin biological interface at the inner face of the plasma membrane, represented in Figure 1(a). The cortex is composed of bundled and cross-linked actin protein filaments that form a dense and overlapping three-dimensional meshwork of a few hundreds of nanometres in thickness [clark2013monitoring]. Myosin-2 molecular motors populate this network and pull on actin filaments by converting chemical into mechanical forces to generate contractile stresses in the meshwork. These internal stresses lead to an effective surface tension over the cell surface and, due to cell curvature, create hydrostatic pressure in the cytoplasm following Young-Laplace’s law. In addition, cortical actin is under constant renewal, through assembly and disassembly of the filaments. This material turnover remodels the whole network in about 30-60 s and releases any elastic stresses accumulated in that period [kelkar2020mechanics]. Being constantly out of thermodynamic equilibrium, the cortex is thus an active system: It behaves as an elastic solid membrane at short time scales (<1 min), but flows like a viscous liquid at longer time scales [prost2015active].
@@
Cortical flows, induced by gradients of cortical tension, are the fundamental mechanism underlying shape change in animal cells [bray1988cortical, mayer2010anisotropies, turlier2014furrow, salbreux2017mechanics]. This is illustrated in Figure 1(b) as follows: A uniform tension in the cortex gives rise to a spherical shape under constant hydrostatic pressure proportional to the surface tension and curvature (here, the inverse of the radius). This is in agreement with normal force balance given by the Young-Laplace law. However, animal cells are able to spatiotemporally localise myosin concentration and motor activity within the cortex, using complex cascades of chemical reactions, known as signalling pathways [Pires-daSilva2003, bement2024patterning]. This can create local imbalances of contractility that, at the time scale of minutes, give rise to Marangoni-like flows, pulling cortical components away from regions of relaxation toward regions of contraction. Moreover, on short time scales, normal force balance implies that gradients of surface tension must necessarily be corresponded with non-uniform cytoplasmic pressures and changes in curvature. While this explains how animal cells exert shape control with the actomyosin cortex, it also reveals that cortex-mediated cell deformation produces a net movement of the cytoplasm [mogilner2018intracellular].
@@
Figure 1. (a) The actomyosin cortex is a thin layer of thickness $t\ll R_{0}$ at the inner face of the plasma membrane, where $R_{0}$ is the typical curvature radius of a cell. It is a network of actin protein filaments and binding proteins, among which myosin-2 molecular motors, who pull on the filaments via a chemical reaction. The network is under constant renewal: polymerisation occurs right under the plasma membrane, while depolymerisation occurs uniformly across the cortex. Adapted from [da2022viscous]. (b) According to Laplace’s law, a uniform tension in the cortex gives rise to a spherical shape under constant hydrostatic pressure. By localising myosin activity, animal cells create gradients of contractility that lead, on long time scales, to tangential cortical flows and, on short time scales, to shape deformation with net cytoplasm movement (non-uniform cytoplasmic pressure). Inspired by [salbreux2012actin], incorporating the appropriate curvature changes associated with increased local membrane tension.
@@
To address this, scholarly efforts combining biophysical experiments and numerical modelling are growing and effectively helping to better understand cortex-cytoplasm interactions [Liao2024, Barnhart2015, bhatnagar2023axis]. The finite element method (FEM) is frequently adopted to approximate the partial differential equations that model these type of systems [torres2022interacting, da2022viscous], given its flexibility to deal with intricate geometries, non-standard boundary conditions and nonlinearities [johnson1987numerical]. Yet, many surface-bulk couplings in living cells involve large distortions of the cell surface. This frames the problem within a classical dilemma in finite element (FE) theory and practice: selecting the description of motion [malvern1969introduction, Donea2004]. This choice constrains the relationship between the deforming body and the computational mesh underlying the FE approximation. As a result, it determines the ability of the FE method to deal with large deformations.
Most, if not all, approaches to describe large distortions with FEs sit between the two main perspectives of motion in continuum mechanics: Lagrangian and Eulerian [Donea2004]. In the Lagrangian viewpoint, the computational grid follows the moving continuum, the grid nodes being permanently identified with the same material points [bennett2006lagrangian]. A natural choice is to use a mesh fitting to the moving boundaries and interfaces in the continuum, i.e., a body-fitted mesh (Figure 2(b)). This perspective is common in nonlinear solid mechanics [bonet1997nonlinear]. While it allows for an easy tracking of free surfaces and interfaces, frequent remeshing is needed to avoid numerical pollution from excessively distorted elements in the mesh [khan2020surface, da2022viscous]. This issue does not appear in the Eulerian viewpoint. In this case, material particles and mesh nodes are decoupled: the computational grid stays fixed, while the continuum moves with respect to the grid (Figure 2(c)). Fixed-grid techniques are predominant in fluid mechanics [anderson1995computational]. Since the mesh no longer fits to moving boundaries and interfaces, tracking their motion becomes more difficult. Besides, convective effects appear, due to the relative motion between the deforming material and the computational grid.
@@
The solid-fluid duality in the mechanics of the actomyosin cortex is rather the rule than the exception in animal cell and tissue interfaces [turlier2014furrow, torres2019modelling, Voigt2019, da2022viscous]. At first glance, there is no clear choice to describe the motion of this type of surfaces and their interactions with surrounding fluids. Hence, it comes as no surprise to see the variety of FE methods in the field reflecting this open-ended question. At one end, there are body-fitted methods [de2021numerical, wittwer2023computational, Ganesan2009, Edelmann2024, hernandez2024bulk], relying on Arbitrary Lagrangian-Eulerian (ALE[Dziuk2013, Mokbel2020, Rangarajan2015] descriptions to cope with the mesh deformation. At the other end, there are fixed-grid methods with diffuse or sharp representations of the moving boundary or interface; they are usually grounded, respectively, on phase-field [kloppe2024phase, teigen2011diffuse] or level-set [Frachon2023] methods. In between, there are hybrid approaches, coupling a body-fitted method for the interface with a fixed-grid one for the bulk phases [Barrett2014, barrett2017finite].
@@
Observe that, of the above works, only body-fitted ALE and diffuse-interface approaches explicitly address applications in cell biology. In particular, as far as the authors know, the most robust and advanced technology in evolving sharp-interface methods remains to be transposed for modelling surface-bulk couplings in animal cells. Working to close this gap could be beneficial in the area for, at least, two reasons: (1) Sharp-interface methods have an apparent edge over body-fitted counterparts when it comes to model cell division, as they do not need to undergo complex numerical surgery for mesh scission [kovacs2024numerical]; (2) Phase-field models can be computationally prohibitive, especially in 3D, as they require small and adaptive mesh resolutions around the moving diffuse-interface [gomez2017computational, Bachini2024].
@@
Given this motivation, the main purpose of this work is to formulate a new sharp-interface FE method for surface-bulk PDEs with dynamic surfaces in fixed computational grids. To approximate the coupled problem, we propose a novel partitioned scheme, in which the surface PDEs are discretised with the trace FEM [Olshanskii2017] and the bulk PDE with the aggregated FEM (AgFEM[badia2018aggregated]. Both methods fall into the class of so-called unfitted, embedded or immersed FE methods [peskin1972flow, Mittal2005]. They are endowed with mathematical properties that ensure numerical stability and robustness, regardless of how the mesh overlaps the moving interface(s) [Hansbo2020, de2023stability], which is an essential requirement in our context. Our contribution distinguishes itself from other unfitted methods in the literature [Frachon2023, Hansbo2016, Olshanskii2022] in two ways: (1) They solve the bulk problem with ghost penalty stabilisation [Burman2015], while we choose AgFEM to avoid locking issues that can appear with standard ghost penalty methods [badia2021linkAgFEM]. (2) We specifically target applications to cortex-cytoplasm fluid interactions in animal cells. For this reason, our model problem consists in a system of surface-bulk viscous flows, whose active driving forces are regulated by the surface transport of a diffusing species. Similar models have only been analysed with body-fitted or diffuse-interface methods [wittwer2023computational, aland2023phase]; in contrast to this work, they do not explicitly solve the surface PDE problem for the viscous flows.
@@
The outline of this paper is as follows. First, we state the problem in Section 2. We introduce minimal notation in Section 2.1 to write the dimensionless 3D governing equations of the problem in Section 2.2. The computational model is detailed in Section 3. We introduce first the embedded geometry setup in Section 3.1. The discretisation in space in Section 3.2 adopts, for the bulk problem, the inf-sup stable pair formed by continuous quadratic velocities and discontinuous linear pressures; for the surface flow problem, continuous linear velocities; and, for the surface transport problem, continuous linear elements. Numerical analysis available for each individual discrete formulation underpin these choices [Badia2018Mixed, jankuhn2021trace]. The discretisation in time considers a simple backward Euler method and the moving surface is modelled with an evolving level-set method (Section 3.3), that leverages high-order quadrature rules on implicit geometries [saye2022high] and extensions of the surface velocity to the whole grid by closest-point projections [saye2014high]. In the numerical experiments of Section 4, we analyse three challenging applications: (1) Self-organised shape emergence, as in [wittwer2023computational], but at large hydrodynamic lengths, (2) relaxation dynamics of 3D bodies and (3) uni- and bi-lateral cell cleavage. Finally, concluding remarks and perspectives are listed in Section 5.
@@ 2. Model problem
We consider the minimal model of self-organization of the cell cortex introduced in [mietke2019minimal]. We restrict our analysis to a single cell. The problem is given by a system of coupled surface-bulk viscous flow equations, where the surface phase represents the cortex and the bulk phase the cytoplasm. Flows are driven by gradients of cortical tension that are regulated by a molecular species diffusing on the cortex, representing the myosin molecular motors. This leads to a system of mixed-dimensional PDEs in the three-dimensional space, where the surface and bulk PDEs are posed in manifolds of dimensionality two and three, respectively.
As we seek to approximate the problem in a fixed three-dimensional grid, it is in our interest to formulate all governing equations in the global Cartesian coordinates. In this way, we do not require local coordinates or a parameterisation for the surface, which facilitates applying the level-set method or other fixed-grid techniques. To achieve this, we will use tangential differential calculus (TDC) to express the surface PDEs in terms of differential operators in Cartesian coordinates, like the bulk PDEs. For the sake of brevity, we will skip most of the mathematical formalism underlying TDC; the interested reader can find it in, e.g., [delfour1996tangential, fries2020unified].
@@ 2.2. Problem statement
Our model problem is an extension of a classical mathematical model for two-phase flow with interfacial tension and viscosity [scriven1960dynamics]. It consists of an active Boussinesq-Scriven surface fluid that encloses a passive viscous bulk fluid, where a diffusing molecular species regulates the driving forces on the surface and creates a mechanochemical feedback.
We focus on cellular processes such as cell division, migration or polarisation, where time scales are long enough to assume the rheology of the cortex is viscous [mayer2010anisotropies, turlier2014furrow]. On the other hand, small characteristic length scales in living cells typically lead to low Reynolds and Womersley numbers ($\mathrm{Re},\ll 1,\,\,\mathrm{\alpha}\ll 1$[mogilner2018intracellular, purcell1977life]. Therefore, neglecting gravitational and inertial forces, balance of momentum on the surface reads
@@
where $\boldsymbol{U}$ denotes the surface velocity field. The first two terms yield the stress resultant owing to viscous effects, with $\mu_{\Gamma}>0$ and $\lambda_{\Gamma}>\mu_{\Gamma}$ the surface shear and dilational viscosities. The last term $\mathbf{N}_{\Gamma}^{\rm act}$ is the contribution from active surface tension. We will show later that this term is the driving force of the system. For the cortex, which is a thin shell of actomyosin material, a dimensional reduction of the 3D constitutive equations to 2D membrane stress resultants yields the relation $\lambda_{\Gamma}=3\mu_{\Gamma}$ under the assumption of cortical incompressibility [da2022viscous]. Additionally, both viscous and active bending contributions have been shown numerically to be negligible [da2022viscous]. We further neglect mechanical contributions from the plasma membrane—specifically, the lipid bilayer’s tangential fluidity and transverse elasticity—including its bending rigidity and passive surface tension. These effects are typically one to several orders of magnitude smaller than the surface stresses generated by the cortex [dimova2014recent, rangamani2022many] and are therefore assumed to have minimal impact on the system’s mechanical behavior.
Equations (2) and (3) imply the key assumption of the cortex being a nonmaterial interface. This means that balance of mass is trivial [bothe2010two], albeit the surface divergence $\mathrm{div}_{\Gamma}\,\boldsymbol{U}$ does not generally vanish. If assuming a material interface, balance of mass can be accounted for in terms of cortical density [torres2019modelling] or thickness [turlier2014furrow]. In this case, material turnover also contributes to the active tension [da2022viscous].
@@
As in [stone1990effects, wittwer2023computational], it is interesting to decompose the force term of the (active) tension in Equation (2) in the local coordinate system of the surface as

### v2

@@ Abstract.
$^{\text{a}}$ Center for Interdisciplinary Research in Biology (CIRB), Collège de France, Université PSL, CNRS, INSERM, Paris, France.
$^{\text{b}}$ Present address: Departament de Mecànica de Fluids (MF), Escola Politècnica Superior d’Enginyeria de Vilanova i la Geltrú (EPSEVG), Universitat Politècnica de Catalunya · BarcelonaTech (UPC), 08800 Vilanova i la Geltrú, Spain.
$^{\text{c}}$ Present address: Center for Integrative Biology (CBI), University of Toulouse, CNRS, Toulouse, France.
@@
E-mails: eric.neiva@upc.edu (EN), herve.turlier@cnrs.fr (HT)
@@ 1. Introduction
Morphogenesis is the biological process by which a cell, tissue or organism takes shape and develops its distinctive form [88]. Researchers in this field seek to understand the underlying genetic, cellular, mechanical, and environmental factors that drive the formation of complex biological structures. The ultimate goal is to unravel the fundamental principles that govern the development of living organisms. Understanding morphogenesis has significant far-reaching implications for various fields, including regenerative medicine [132], cancer research [158] or environmental and ecological studies [156].
@@
This work centres upon the morphogenesis of single animal cells. Animal cells maintain their shape primarily through an effective surface tension, adopting shapes akin to soap bubbles [72]. The main contributor to this effective surface tension in animal cells is the actomyosin cortex [38]: a thin biological interface at the inner face of the plasma membrane, represented in Figure 1(a). The cortex is composed of bundled and cross-linked actin protein filaments that form a dense and overlapping three-dimensional meshwork of a few hundreds of nanometres in thickness [37]. Myosin-2 molecular motors populate this network and pull on actin filaments by converting chemical into mechanical forces to generate contractile stresses in the meshwork. These internal stresses lead to an effective surface tension over the cell surface and, due to cell curvature, create hydrostatic pressure in the cytoplasm following Young-Laplace’s law. In addition, cortical actin is under constant renewal, through assembly and disassembly of the filaments. This material turnover remodels the whole network in about 30-60 s and releases any elastic stresses accumulated in that period [83]. Being constantly out of thermodynamic equilibrium, the cortex is thus an active system: It behaves as an elastic solid membrane at short time scales (<1 min), but flows like a viscous liquid at longer time scales [119].
@@
Cortical flows, induced by gradients of cortical tension, are the fundamental mechanism underlying shape change in animal cells [28, 97, 148, 130]. This is illustrated in Figure 1(b) as follows: A uniform tension in the cortex gives rise to a spherical shape under constant hydrostatic pressure proportional to the surface tension and curvature (here, the inverse of the radius). This is in agreement with normal force balance given by the Young-Laplace law. However, animal cells are able to spatiotemporally localise myosin concentration and motor activity within the cortex, using complex cascades of chemical reactions, known as signalling pathways [117, 18]. This can create local imbalances of contractility that, at the time scale of minutes, give rise to Marangoni-like flows, pulling cortical components away from regions of relaxation toward regions of contraction. Moreover, on short time scales, normal force balance implies that gradients of surface tension must necessarily be corresponded with non-uniform cytoplasmic pressures and changes in curvature. While this explains how animal cells exert shape control with the actomyosin cortex, it also reveals that cortex-mediated cell deformation produces a net movement of the cytoplasm [103].
@@
Figure 1. (a) The actomyosin cortex is a thin layer of thickness $t\ll R_{0}$ at the inner face of the plasma membrane, where $R_{0}$ is the typical curvature radius of a cell. It is a network of actin protein filaments and binding proteins, among which myosin-2 molecular motors, who pull on the filaments via a chemical reaction. The network is under constant renewal: polymerisation occurs right under the plasma membrane, while depolymerisation occurs uniformly across the cortex. Adapted from [40]. (b) According to Laplace’s law, a uniform tension in the cortex gives rise to a spherical shape under constant hydrostatic pressure. By localising myosin activity, animal cells create gradients of contractility that lead, on long time scales, to tangential cortical flows and, on short time scales, to shape deformation with net cytoplasm movement (non-uniform cytoplasmic pressure). Inspired by [128], incorporating the appropriate curvature changes associated with increased local membrane tension.
@@
To address this, scholarly efforts combining biophysical experiments and numerical modelling are growing and effectively helping to better understand cortex-cytoplasm interactions [91, 14, 23]. The finite element method (FEM) is frequently adopted to approximate the partial differential equations that model these type of systems [145, 40], given its flexibility to deal with intricate geometries, non-standard boundary conditions and nonlinearities [82]. Yet, many surface-bulk couplings in living cells involve large distortions of the cell surface. This frames the problem within a classical dilemma in finite element (FE) theory and practice: selecting the description of motion [95, 49]. This choice constrains the relationship between the deforming body and the computational mesh underlying the FE approximation. As a result, it determines the ability of the FE method to deal with large deformations.
Most, if not all, approaches to describe large distortions with FEs sit between the two main perspectives of motion in continuum mechanics: Lagrangian and Eulerian [49]. In the Lagrangian viewpoint, the computational grid follows the moving continuum, the grid nodes being permanently identified with the same material points [19]. A natural choice is to use a mesh fitting to the moving boundaries and interfaces in the continuum, i.e., a body-fitted mesh (Figure 2(b)). This perspective is common in nonlinear solid mechanics [25]. While it allows for an easy tracking of free surfaces and interfaces, frequent remeshing is needed to avoid numerical pollution from excessively distorted elements in the mesh [84, 40]. This issue does not appear in the Eulerian viewpoint. In this case, material particles and mesh nodes are decoupled: the computational grid stays fixed, while the continuum moves with respect to the grid (Figure 2(c)). Fixed-grid techniques are predominant in fluid mechanics [5]. Since the mesh no longer fits to moving boundaries and interfaces, tracking their motion becomes more difficult. Besides, convective effects appear, due to the relative motion between the deforming material and the computational grid.
@@
The solid-fluid duality in the mechanics of the actomyosin cortex is rather the rule than the exception in animal cell and tissue interfaces [148, 146, 153, 40]. At first glance, there is no clear choice to describe the motion of this type of surfaces and their interactions with surrounding fluids. Hence, it comes as no surprise to see the variety of FE methods in the field reflecting this open-ended question. At one end, there are body-fitted methods [43, 157, 59, 51, 73], relying on Arbitrary Lagrangian-Eulerian (ALE[50, 104, 122] descriptions to cope with the mesh deformation. At the other end, there are fixed-grid methods with diffuse or sharp representations of the moving boundary or interface; they are usually grounded, respectively, on phase-field [85, 143] or level-set [56] methods. In between, there are hybrid approaches, coupling a body-fitted method for the interface with a fixed-grid one for the bulk phases [16, 15].
@@
Observe that, of the above works, only body-fitted ALE and diffuse-interface approaches explicitly address applications in cell biology. In particular, as far as the authors know, the most robust and advanced technology in evolving sharp-interface methods remains to be transposed for modelling surface-bulk couplings in animal cells. Working to close this gap could be beneficial in the area for, at least, two reasons: (1) Sharp-interface methods have an apparent edge over body-fitted counterparts when it comes to model cell division, as they do not need to undergo complex numerical surgery for mesh scission [86]; (2) Phase-field models can be computationally prohibitive, especially in 3D, as they require small and adaptive mesh resolutions around the moving diffuse-interface [62, 6].
@@
Given this motivation, the main purpose of this work is to formulate a new sharp-interface FE method for surface-bulk PDEs with dynamic surfaces in fixed computational grids. To approximate the coupled problem, we propose a novel partitioned scheme, in which the surface PDEs are discretised with the trace FEM [113] and the bulk PDE with the aggregated FEM (AgFEM[9]. Both methods fall into the class of so-called unfitted, embedded or immersed FE methods [116, 100]. They are endowed with mathematical properties that ensure numerical stability and robustness, regardless of how the mesh overlaps the moving interface(s) [68, 44], which is an essential requirement in our context. Our contribution distinguishes itself from other unfitted methods in the literature [56, 69, 114] in two ways: (1) They solve the bulk problem with ghost penalty stabilisation [30], while we choose AgFEM to avoid locking issues that can appear with standard ghost penalty methods [8]. (2) We specifically target applications to cortex-cytoplasm fluid interactions in animal cells. For this reason, our model problem consists in a system of surface-bulk viscous flows, whose active driving forces are regulated by the surface transport of a diffusing species. Similar models have only been analysed with body-fitted or diffuse-interface methods [157, 2]; in contrast to this work, they do not explicitly solve the surface PDE problem for the viscous flows.
@@
The outline of this paper is as follows. First, we state the problem in Section 2. We introduce minimal notation in Section 2.1 to write the dimensionless 3D governing equations of the problem in Section 2.2. The computational model is detailed in Section 3. We introduce first the embedded geometry setup in Section 3.1. The discretisation in space in Section 3.2 adopts, for the bulk problem, the inf-sup stable pair formed by continuous quadratic velocities and discontinuous linear pressures; for the surface flow problem, continuous linear velocities; and, for the surface transport problem, continuous linear elements. Numerical analysis available for each individual discrete formulation underpin these choices [7, 81]. The discretisation in time considers a simple backward Euler method and the moving surface is modelled with an evolving level-set method (Section 3.3), that leverages high-order quadrature rules on implicit geometries [134] and extensions of the surface velocity to the whole grid by closest-point projections [133]. In the numerical experiments of Section 4, we analyse three challenging applications: (1) Self-organised shape emergence, as in [157], but at large hydrodynamic lengths, (2) relaxation dynamics of 3D bodies and (3) uni- and bi-lateral cell cleavage. Finally, concluding remarks and perspectives are listed in Section 5.
@@ 2. Model problem
We consider the minimal model of self-organization of the cell cortex introduced in [98]. We restrict our analysis to a single cell. The problem is given by a system of coupled surface-bulk viscous flow equations, where the surface phase represents the cortex and the bulk phase the cytoplasm. Flows are driven by gradients of cortical tension that are regulated by a molecular species diffusing on the cortex, representing the myosin molecular motors. This leads to a system of mixed-dimensional PDEs in the three-dimensional space, where the surface and bulk PDEs are posed in manifolds of dimensionality two and three, respectively.
As we seek to approximate the problem in a fixed three-dimensional grid, it is in our interest to formulate all governing equations in the global Cartesian coordinates. In this way, we do not require local coordinates or a parameterisation for the surface, which facilitates applying the level-set method or other fixed-grid techniques. To achieve this, we will use tangential differential calculus (TDC) to express the surface PDEs in terms of differential operators in Cartesian coordinates, like the bulk PDEs. For the sake of brevity, we will skip most of the mathematical formalism underlying TDC; the interested reader can find it in, e.g., [46, 58].
@@ 2.2. Problem statement
Our model problem is an extension of a classical mathematical model for two-phase flow with interfacial tension and viscosity [138]. It consists of an active Boussinesq-Scriven surface fluid that encloses a passive viscous bulk fluid, where a diffusing molecular species regulates the driving forces on the surface and creates a mechanochemical feedback.
We focus on cellular processes such as cell division, migration or polarisation, where time scales are long enough to assume the rheology of the cortex is viscous [97, 148]. On the other hand, small characteristic length scales in living cells typically lead to low Reynolds and Womersley numbers ($\mathrm{Re},\ll 1,\,\,\mathrm{\alpha}\ll 1$[103, 120]. Therefore, neglecting gravitational and inertial forces, balance of momentum on the surface reads
@@
where $\boldsymbol{U}$ denotes the surface velocity field. The first two terms yield the stress resultant owing to viscous effects, with $\mu_{\Gamma}>0$ and $\lambda_{\Gamma}>\mu_{\Gamma}$ the surface shear and dilational viscosities. The last term $\mathbf{N}_{\Gamma}^{\rm act}$ is the contribution from active surface tension. We will show later that this term is the driving force of the system. For the cortex, which is a thin shell of actomyosin material, a dimensional reduction of the 3D constitutive equations to 2D membrane stress resultants yields the relation $\lambda_{\Gamma}=3\mu_{\Gamma}$ under the assumption of cortical incompressibility [40]. Additionally, both viscous and active bending contributions have been shown numerically to be negligible [40]. We further neglect mechanical contributions from the plasma membrane—specifically, the lipid bilayer’s tangential fluidity and transverse elasticity—including its bending rigidity and passive surface tension. These effects are typically one to several orders of magnitude smaller than the surface stresses generated by the cortex [48, 121] and are therefore assumed to have minimal impact on the system’s mechanical behavior.
Equations (2) and (3) imply the key assumption of the cortex being a nonmaterial interface. This means that balance of mass is trivial [26], albeit the surface divergence $\mathrm{div}_{\Gamma}\,\boldsymbol{U}$ does not generally vanish. If assuming a material interface, balance of mass can be accounted for in terms of cortical density [146] or thickness [148]. In this case, material turnover also contributes to the active tension [40].
@@
As in [141, 157], it is interesting to decompose the force term of the (active) tension in Equation (2) in the local coordinate system of the surface as